# RoboMemory: A Brain-inspired Multi-memory Agentic Framework for Lifelong Learning in Physical Embodied Systems

### v3

@@ A.1 Evaluate on EB-Habitat
Navigate to object/navigation point find(obj) navigate(point) navigate_to(point)
@@
Appendix B Additional Environment settings
B.1 EB-ALFRED and EB-Habitat
We adopt the same environment parameters as in EmbodiedBench. The maximum steps per task are set to 30, with image inputs of size 500 $\times$ 500. The temporal memory buffer length is set to 3. However, we modify the action formats of EB-ALFRED and EB-Habitat to better simulate real-world scenarios. Specifically, we define different action APIs (Python functions), where each action takes an object parameter indicating its target. We extract all possible objects from the environment as inputs to the Agent. The Agent must select appropriate actions and object parameters based on task requirements. Compared to the original interaction method in EmbodiedBench (which enumerates all possible actions, including both action names and target objects, and requires the Agent to choose), our approach offers greater flexibility. The detailed action APIs are presented in Table 3.
B.2 Real-world experiments
We construct a common kitchen scenario to evaluate the RoboMemory framework’s lifelong learning capabilities in real-world settings. Using Mobile ALOHA [14] as our physical robotic platform, we design three categories of tasks: (1) Pick up & put down: The agent must locate a specified object among all possible positions and place it at a designated location. This task tests the model’s basic object-searching and planning abilities. (2) Pick up, operate & put down: Building upon the first task, the agent must additionally perform operations such as heating or cleaning the object. This task requires longer-term planning, which is crucial in embodied environments. (3) Pick up, gather & put down: The agent must place specified objects into a movable container and then move the container to a target location. This task evaluates the agent’s understanding of object relationships, requiring it to remember the positions of at least two objects (the container and the target item) and their spatial relationship.
@@ B.2 Real-world experiments
In our experiments, we set the maximum steps per task to 15. We also provide an API for actively terminating tasks. Since real-world environments lack direct success/failure feedback, RoboMemory must autonomously determine task completion. To prevent excessively long task execution, we enforce termination after 15 steps if no success is achieved. A single main camera (640 $\times$ 480 resolution) records video during action execution as input for RoboMemory’s higher-level processing.
@@
B.3 Training Details of Low-Level Executer
We use the $\pi_{0}$ model as our foundation model. We collected 1,040 data samples over 10 types of tasks for fine-tuning. We use LoRA [17] fine-tuning to save resources during fine-tuning. The specific fine-tuning parameters and action types are given in Table 4. For tasks involving both pick-up and place actions, we split these tasks into separate pick-up and place actions. These are then treated as two distinct data samples during training. The separation of pick-up and place action allows the VLA to carry an object in its hand. For training, we used a server with six A100-80GB GPUs. The total training time was 12 hours.
@@ B.3 Training Details of Low-Level Executer
Appendix C Proof of Dynamic Spatial Memory Update Algorithm
Theorem 1 (Upper Bound on K-hop Node Extraction in Directed Graphs).
Let $G=(V,E)$ be a finite directed graph with maximum out-degree $D\geq 1$, and let $\mathcal{S}\subseteq V$ be a set of $M$ source nodes. Define the K-hop neighborhood $\mathcal{N}_{K}(s)$ of a node $s\in\mathcal{S}$ as the set of nodes reachable from $s$ via directed paths of length at most $K$. Then the total number of distinct nodes in the union of all K-hop neighborhoods,
$\mathcal{N}_{K}(\mathcal{S})=\bigcup_{s\in\mathcal{S}}\mathcal{N}_{K}(s),$
Satisfies the following upper bound:
Proof.
For any node $s\in\mathcal{S}$, the number of distinct nodes reachable from $s$ within $i$ hops is at most $D^{i}$, assuming the worst-case scenario where each node encountered has the maximum out-degree $D$, and all neighbors are distinct and non-overlapping.
Thus, the size of the K-hop neighborhood of a single node satisfies:
Since there are $M$ such source nodes and assuming no overlaps between their K-hop neighborhoods (worst case), the union size satisfies:
$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot|\mathcal{N}_{K}(s)|.$
Substituting the bound on $|\mathcal{N}_{K}(s)|$ gives the result. ∎
Theorem 2 (Upper Bound for K-hop Node Extraction in Normalized Directed Graphs).
Let $G=(V,E)$ be a finite directed graph with $|V|=n$ nodes. Assume the maximum out-degree is at most $D_{\max}=Dn$, and the maximum in-degree is at most $N_{\max}=Nn$, where $D,N\in(0,1]$ are constants. Let $\mathcal{S}\subseteq V$ be a set of $M$ source nodes. Define $\mathcal{N}_{K}(\mathcal{S})$ as the union of all nodes reachable from $\mathcal{S}$ via paths of length at most $K$, using only outgoing edges. Then the number of extracted nodes satisfies:
In particular, when $Dn\gg 1$, we have the approximation:
$|\mathcal{N}_{K}(\mathcal{S})|\lessapprox M\cdot(Dn)^{K}.$
Proof.
For each node $s\in\mathcal{S}$, the maximum number of reachable nodes within $i$-hops is at most $(Dn)^{i}$ under the assumption of maximum out-degree and no overlap.
Summing over hops from 0 to $K$, we get for each root:
$|\mathcal{N}_{K}(s)|\leq\sum_{i=0}^{K}(Dn)^{i}=\frac{(Dn)^{K+1}-1}{Dn-1}.$
Assuming no overlap among the $M$ source node expansions (worst case), we have:
$|\mathcal{N}_{K}(\mathcal{S})|\leq M\cdot\frac{(Dn)^{K+1}-1}{Dn-1}.$
Since the total number of nodes in the graph is $n$, this quantity is also trivially bounded above by $n$, yielding the result. ∎
Appendix D Supplementary Examples for Qualitative Analysis
Figure 7: Visualization of Spatial Memory’s dynamic update process.
D.1 Example of Dynamic Spatial Memory Update process
In RoboMemory’s Spatial Memory, the KG is dynamically constructed during environment exploration. As illustrated in Fig. 7, we demonstrate the progressive expansion of the KG in Spatial Memory as the agent navigates through the environment. The figure indicates a continuous growth in the number of both nodes and edges of the KG as exploration progresses.
D.2 Real World
Figure 8: Case that a task is failed but the experience can help RoboMemory to succeed in the next try.
In Fig. 8, we demonstrate an example of RoboMemory learning through trial and error in a real-world environment. Our task is “place a banana into the oven.” This task required RoboMemory to complete the objectives of finding the banana, picking it up, and transporting it to the oven. We observed that RoboMemory became stuck in an infinite loop during the first attempt. The banana was randomly placed on the “kitchen counter,” but RoboMemory overlooked this navigation target and remained trapped exploring other navigation targets instead.
@@ D.2 Real World
We also provide an example that completes the task in the first attempt. The example is shown in Fig. 9. This example demonstrates that the RoboMemory has the ability to handle some relatively complex tasks in the real world. The task in this example is “Place a box of gum into the basket and put the basket on the kitchen counter”. Because two objects in different positions are involved in this task, RoboMemory has to memorize the position of at least one object to achieve the goal. With the help of the spatial memory, RoboMemory completes the task successfully.
D.3 EB-ALFRED
@@ D.3.1 Successful example
We select a successful example to show how RoboMemory performed in the EB-ALFRED environment. The example trajectory is shown in Fig. 10.
@@ D.3.2 Failed example
We demonstrate a representative example of the Critical Error. The example trajectory is shown in Fig. 11. In this example, the task involves slicing and heating a tomato and moving the heated tomato slice to the trash can. Initially, RoboMemory successfully sliced the tomato with a knife. But when the planner plans the whole sequence, it forgets to drop the knife before picking up the tomato (this is necessary because in EB-ALFRED, the robot can only hold one object at a time). The critic and the planner should notice this situation and ask the critic to replan, as RoboMemory failed to pick up a tomato slice. However, the critic module ignores this issue, and thus, after it heats the knife instead of a tomato slice, it stacks in an infinite loop.
@@
Besides, we provide another example demonstrating a representative failure caused by inaccurate action planning. The example trajectory is shown in Fig. 12. In the trajectory, RoboMemory is asked to place two CDs into the drawer. However, at step 6, the robot failed to select correct CD object. In this experiment, RoboMemory has already put CD_2 into the drawer, but it keeps picking up CD_2 even though the memory has clearly indicated that CD_2 has already been put down. So we classify this as inaccurate action error. This indicates that the planner failed to comprehensively integrate information from both the memory and information-gathering modules, resulting in inaccurate action planning.

### v4

@@ D.2 Real-world experiments
Navigation find(obj) navigate(point) navigate_to(point)
@@
We construct a common kitchen scenario to evaluate the RoboMemory framework’s interactive environmental learning capabilities in real-world settings. Using Mobile ALOHA [15] as our physical robotic platform, we design three categories of tasks: (1) Pick up & put down: The agent must locate a specified object among all possible positions and place it at a designated location. This task tests the model’s basic object-searching and planning abilities. (2) Pick up, operate & put down: Building upon the first task, the agent must additionally perform operations such as heating or cleaning the object. This task requires longer-term planning, which is crucial in embodied environments. (3) Pick up, gather & put down: The agent must place specified objects into a movable container and then move the container to a target location. This task evaluates the agent’s understanding of object relationships, requiring it to remember the positions of at least two objects (the container and the target item) and their spatial relationship. For each type of task, we design 5 tasks. So our experiments include 15 long-term real-world tasks.
@@
In our experiments, we set the maximum steps per task to 25. We also provide an API for actively terminating tasks. Since real-world environments lack direct success/failure feedback, RoboMemory must autonomously determine task completion. To prevent excessively long task execution, we enforce termination after 25 steps if no success is achieved. A single main camera (640 $\times$ 480 resolution) records video during action execution as input for RoboMemory’s higher-level processing.
@@
D.3 Training Details of Low-Level Executor
We use the $\pi_{0}$ model as our foundation model. We collected 1,040 data samples over 10 types of tasks for fine-tuning. We use LoRA fine-tuning to save resources during fine-tuning. The specific fine-tuning parameters and action types are given in Table 5. For tasks involving both pick-up and place actions, we split these tasks into separate pick-up and place actions. These are then treated as two distinct data samples during training. The separation of pick-up and place action allows the VLA to carry an object in its hand. For training, we used a server with six A100-80GB GPUs. The total training time was 12 hours.
@@ D.3 Training Details of Low-Level Executor
D.4 Hyperparameters of RoboMemory
In this section, we describe the hyperparameter settings for the upper brain of RoboMemory: the information preprocessor and the Comprehensive Embodied Memory. Importantly, we use a unified set of hyperparameters across all experimental settings, including EB-ALFRED, EB-Habitat, and real-world deployments.
The information preprocessor consists of two components: a step summarizer and a query generator. Given multimodal inputs at each step, the step summarizer produces a single natural language description, while the query generator concurrently formulates 454-5 distinct natural language queries.
The Comprehensive Embodied Memory integrates four memory modules: Temporal, Spatial, Semantic, and Episodic Memory. The Temporal Memory is implemented as a fixed-size buffer with a maximum capacity of $4$ entries. For Spatial Memory, during similarity-based retrieval, we first identify the top $N=3$ most relevant vertices and then perform a K-hop graph traversal with $K=2$. The Episodic Memory retrieves the top $N=5$ most relevant past experiences for each query. The Semantic Memory maintains hierarchical summaries at both the action and task levels; during retrieval, it returns $N_{s}=2$ action-level and $N_{t}=2$ task-level summaries. Furthermore, memory updates (e.g., insertion, modification, or deletion) are applied only to the top $N_{\text{update}}=10$ most relevant entries in the Semantic Memory to ensure efficiency and coherence.
Appendix E Supplementary Examples for Qualitative Analysis
E.1 Real World
Figure 10: Case that a task is failed, but the experience can help RoboMemory to succeed in the next try.
In Figure 10, we demonstrate an example of RoboMemory learning through trial and error in a real-world environment. Our task is “place a banana into the oven.” This task required RoboMemory to complete the objectives of finding the banana, picking it up, and transporting it to the oven. We observed that RoboMemory became stuck in an infinite loop during the first attempt. The banana was randomly placed on the “kitchen counter,” but RoboMemory overlooked this navigation target and remained trapped, exploring other navigation targets instead.
@@ E.1 Real World
We also provide an example that completes the task in the first attempt. The example is shown in Figure 11. This example demonstrates that the RoboMemory has the ability to handle some relatively complex tasks in the real world. The task in this example is “Place a box of gum into the basket and put the basket on the kitchen counter”. Because two objects in different positions are involved in this task, RoboMemory has to memorize the position of at least one object to achieve the goal. With the help of the spatial memory, RoboMemory completes the task successfully.
E.2 EB-ALFRED
@@ E.2.1 Successful example
We select a successful example to show how RoboMemory performed in the EB-ALFRED environment. The example trajectory is shown in Figure 12.
@@ E.2.2 Failed example
We demonstrate a representative example of the Critical Error. The example trajectory is shown in Figure 13. In this example, the task involves slicing and heating a tomato and moving the heated tomato slice to the trash can. Initially, RoboMemory successfully sliced the tomato with a knife. But when the planner plans the whole sequence, it forgets to drop the knife before picking up the tomato (this is necessary because in EB-ALFRED, the robot can only hold one object at a time). The critic and the planner should notice this situation and ask the critic to replan, as RoboMemory failed to pick up a tomato slice. However, the critic module ignores this issue, and thus, after it heats the knife instead of a tomato slice, it stacks in an infinite loop.
@@
Besides, we provide another example demonstrating a representative failure caused by inaccurate action planning. The example trajectory is shown in Figure 14. In the trajectory, RoboMemory is asked to place two CDs into the drawer. However, at step 6, the robot failed to select correct CD object. In this experiment, RoboMemory has already put CD_2 into the drawer, but it keeps picking up CD_2 even though the memory has clearly indicated that CD_2 has already been put down. So we classify this as inaccurate action error. This indicates that the planner failed to comprehensively integrate information from both the memory and information-gathering modules, resulting in inaccurate action planning.